i dont know what category this goes in but:
the sum of twice a number and three times its square is 261. Find the number.

Answer by nerdybill(7384)

the sum of twice a number and three times its square is 261. Find the number.
.
Let n = the number
then
2n + 3n^2 = 261
3n^2 + 2n = 261
3n^2 + 2n - 261 = 0
factoring the left:
(3n+29)(n-9) = 0
n = {-29/3, 9}
answer: 9